OpenStudy (anonymous):

find rectangular coordinates of (-7, 5 pi/2)

OpenStudy (anonymous):

\[(-7,\frac{5\pi}{2}) = (r,\theta)\]\[x=r \cos \theta\]\[y=r \sin \theta\]

OpenStudy (anonymous):

I got rcosx=0, and rsinx=-7

OpenStudy (anonymous):

is this correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the rectangular coordinate is (x,y), its done

OpenStudy (anonymous):

oh sick I think I understand them now thanks